# Forget Vectors at Play: Universal Input Perturbations Driving Machine Unlearning in Image Classification

### v2

@@ 6.2 Experiment Results
Table 1: Performance overview of various MU methods for image classification under two unlearning scenarios on CIFAR-10 and ImageNet-10 using ResNet-18 and VGG-16, respectively. Results are reported in the format $a_{\pm b}$, where $a$ is the mean and $b$ denotes standard deviation $b$ over $10$ independent trials. The performance gap against Retrain is indicated in ($\bullet$), where a lower value is better. $\uparrow$ (or $\downarrow$) indicates that a higher (or lower) value is better. The best performance for each metric is highlighted in green, while the second-best performance is highlighted in red.
[…]
Class-wise Forgetting, CIFAR-10 Class-wise Forgetting, ImageNet-10
Retrain $100.00_{\pm{0.00}}$($0.00$) $100.00_{\pm{0.00}}$($0.00$) $99.91_{\pm{0.03}}$($0.00$) $94.92_{\pm{0.15}}$($0.00$) $0.00$ $100.00_{\pm{0.00}}$($0.00$) $100.00_{\pm{0.00}}$($0.00$) $99.66_{\pm{0.16}}$($0.00$) $97.11_{\pm{0.82}}$($0.00$) $0.00$
FT $5.27_{\pm{0.73}}$($94.73$) $51.49_{\pm{5.07}}$($48.51$) $100.0_{\pm{0.0}}$($0.09$) $95.03_{\pm{0.07}}$($0.11$) $35.86$ $39.66_{\pm{4.73}}$($60.34$) $55.76_{\pm{7.26}}$($44.24$) $99.78_{\pm{0.03}}$($0.13$) $97.27_{\pm{0.35}}$($2.35$) $26.77$
RL $18.87_{\pm{7.34}}$($81.13$) $98.94_{\pm{0.79}}$($1.06$) $99.98_{\pm{0.0}}$($0.07$) $94.51_{\pm{0.12}}$($0.41$) $20.67$ $76.58_{\pm{11.64}}$($23.42$) $46.04_{\pm{33.71}}$($53.96$) $99.28_{\pm{0.2}}$($0.63$) $96.91_{\pm{0.55}}$($1.99$) $20.00$
GA $71.45_{\pm{0.35}}$($28.55$) $81.7_{\pm{0.22}}$($18.30$) $98.62_{\pm{0.04}}$($1.29$) $92.34_{\pm{0.02}}$($2.58$) $12.68$ $46.61_{\pm{6.11}}$($53.39$) $49.15_{\pm{9.36}}$($50.85$) $99.35_{\pm{0.11}}$($0.56$) $95.6_{\pm{0.22}}$($0.68$) $26.37$
NegGrad+ $91.78_{\pm{14.66}}$($8.22$) $95.81_{\pm{7.28}}$($4.19$) $98.35_{\pm{1.22}}$($1.56$) $92.62_{\pm{1.34}}$($2.3$) $4.07$ $49.56_{\pm{34.87}}$($50.44$) $64.27_{\pm{27.33}}$($35.73$) $99.08_{\pm{1.18}}$($0.58$) $96.47_{\pm{1.25}}$($0.64$) $21.84$
SalUn $96.35_{\pm{2.14}}$($3.65$) $98.64_{\pm{0.03}}$($1.36$) $98.75_{\pm{0.18}}$($1.16$) $92.34_{\pm{1.54}}$($2.58$) $2.19$ $95.13_{\pm{1.79}}$($4.87$) $97.24_{\pm{0.17}}$($2.76$) $96.33_{\pm{0.25}}$($3.33$) $96.18_{\pm{1.10}}$($0.93$) $2.97$
SCRUB $93.45_{\pm{2.33}}$($6.55$) $96.38_{\pm{1.72}}$($3.62$) $99.95_{\pm{0.0}}$($0.04$) $94.56_{\pm{0.07}}$($0.36$) $2.64$ $99.12_{\pm{0.14}}$($0.88$) $98.01_{\pm{0.51}}$($1.99$) $99.75_{\pm{0.03}}$($0.09$) $97.24_{\pm{0.16}}$($0.13$) $0.77$
Ours $97.88_{\pm{0.27}}$($2.12$) $99.60_{\pm{0.15}}$($0.40$) $97.25_{\pm{0.24}}$($2.66$) $90.90_{\pm{0.32}}$($4.02$) $2.30$ $87.23_{\pm{6.55}}$($12.77$) $91.41_{\pm{5.9}}$($8.59$) $94.77_{\pm{1.16}}$($5.14$) $94.04_{\pm{1.29}}$($0.88$) $6.85$
Random Data Forgetting, CIFAR-10 Random Data Forgetting, ImageNet-10
Retrain $5.50_{\pm{0.16}}$($0.00$) $11.57_{\pm{0.47}}$($0.00$) $99.88_{\pm{0.05}}$($0.00$) $94.24_{\pm{0.19}}$($0.00$) $0.00$ $4.05_{\pm{0.45}}$($0.00$) $6.60_{\pm{1.07}}$($0.00$) $99.48_{\pm{0.07}}$($0.00$) $96.33_{\pm{0.38}}$($0.00$) $0.00$
FT $0.03_{\pm{0.03}}$($5.47$) $0.75_{\pm{0.09}}$($10.82$) $99.98_{\pm{0.02}}$($0.10$) $94.45_{\pm{0.14}}$($0.21$) $4.15$ $1.35_{\pm{0.32}}$($2.70$) $4.67_{\pm{1.61}}$($1.93$) $99.36_{\pm{0.28}}$($0.12$) $96.54_{\pm{0.59}}$($0.21$) $1.24$
RL $0.52_{\pm{0.24}}$($4.98$) $3.13_{\pm{0.55}}$($8.44$) $99.85_{\pm{0.07}}$($0.03$) $93.88_{\pm{0.20}}$($0.36$) $3.45$ $2.96_{\pm{0.42}}$($1.09$) $12.85_{\pm{4.25}}$($6.25$) $99.19_{\pm{0.16}}$($0.29$) $95.50_{\pm{0.9}}$($0.83$) $2.12$
GA $1.56_{\pm{3.08}}$($3.94$) $2.88_{\pm{3.44}}$($8.69$) $98.67_{\pm{2.74}}$($1.21$) $92.84_{\pm{2.59}}$($1.40$) $3.81$ $0.18_{\pm{0.04}}$($3.87$) $2.97_{\pm{1.51}}$($3.63$) $99.86_{\pm{0.01}}$($0.38$) $97.47_{\pm{0.09}}$($1.14$) $2.23$
NegGrad+ $0.97_{\pm{1.08}}$($4.53$) $2.74_{\pm{2.16}}$($8.83$) $99.42_{\pm{0.87}}$($0.46$) $93.38_{\pm{1.13}}$($0.86$) $3.67$ $0.60_{\pm{0.44}}$($3.45$) $3.90_{\pm{2.22}}$($2.70$) $99.68_{\pm{0.28}}$($0.20$) $97.12_{\pm{0.45}}$($0.79$) $1.79$
SalUn $1.73_{\pm{0.25}}$($3.77$) $6.25_{\pm{1.21}}$($5.32$) $99.24_{\pm{0.09}}$($0.64$) $91.03_{\pm{0.14}}$($3.21$) $3.23$ $1.15_{\pm{0.40}}$($2.90$) $3.56_{\pm{1.12}}$($3.04$) $98.97_{\pm{1.56}}$($0.51$) $95.42_{\pm{2.10}}$($0.91$) $1.84$
SCRUB $0.61_{\pm{0.31}}$($4.89$) $3.69_{\pm{0.54}}$($7.88$) $99.76_{\pm{0.18}}$($0.12$) $93.91_{\pm{0.19}}$($0.33$) $3.31$ $0.18_{\pm{0.08}}$($3.87$) $2.80_{\pm{1.33}}$($3.80$) $99.88_{\pm{0.03}}$($0.40$) $97.36_{\pm{0.15}}$($1.03$) $2.28$
Ours $2.61_{\pm{0.49}}$($2.89$) $8.26_{\pm{1.17}}$($3.00$) $97.33_{\pm{0.47}}$($2.55$) $90.97_{\pm{0.38}}$($3.27$) $2.92$ $2.27_{\pm{0.50}}$($1.78$) $6.13_{\pm{1.40}}$($0.47$) $98.29_{\pm{0.32}}$($1.19$) $95.82_{\pm{0.48}}$($0.51$) $0.99$
@@
In Tab. 1, we compare the performance of our forget vector approach with other model-based MU methods across the metrics: UA, RA, TA, MIA-Efficacy, and Avg. Gap vs. Retrain. We highlight two key observations below. First, in terms of unlearning effectiveness (UA and MIA-Efficacy), the data perturbation-based forget vector demonstrates highly competitive performance compared to model update-based MU baselines, typically ranking among the top two methods with the smallest performance gap relative to Retrain. The advantage of the forget vector is particularly evident in MIA-Efficacy, where it achieves the closest results to Retrain. Second, in terms of model utility post-unlearning (RA and TA), the forget vector generally leads to a larger performance drop than other methods. This is not surprising, as the forget vector introduces data perturbations to the retain and testing tests at evaluation while keeping the model unchanged, even though it is designed to preserve model utility in (6). Nonetheless, a certain RA or TA loss may still occur. However, considering the gain in unlearning effectiveness, the Avg. Gap with Retrain shows that the forget vector remains competitive, ranking among the top two unlearning methods. Furthermore, we remark that the utility loss caused by the forget vector could be mitigated by refraining from applying it to utility datasets, as its primary role is to proactively enforce forgetting on the forget set.
Table 2: Unlearning accuracy (%) of forget vector transferred to unseen forget sets under 3 scenarios on (CIFAR-10, ResNet-18).
@@
Conventionally, unlearning effectiveness is typically measured on the original forget set ($\mathcal{D}_{\mathrm{f}}$). However, with the data perturbation-based forget vector, it is also interesting to investigate its unlearning transferability when applied to a new, previously unseen forget set (denoted as $\mathcal{D_{\mathrm{f}}}^{\prime}$) that share similarities with $\mathcal{D}_{\mathrm{f}}$ and are equally appropriate for unlearning. In the context of class-wise forgetting, we consider $\mathcal{D}_{\mathrm{f}}^{\prime}$ using the testing data from the class targeted for unlearning. In the context of data forgetting, we obtain $\mathcal{D}_{\mathrm{f}}^{\prime}$ by applying the data corruption operation GN1 and ET1 used in Fig. 2 to perturb $\mathcal{D}_{\mathrm{f}}$. Tab. 2 shows the UA of the forget vector applied to the unseen forget set $\mathcal{D}_{\mathrm{f}}^{\prime}$. As observed, the unlearning performance of the forget vector remains effective when evaluated on $\mathcal{D}_{\mathrm{f}}^{\prime}$. Among the model update-based unlearning baselines, current SOTA methods such as SCRUB (Kurmanji et al., 2024b) and SalUn (Fan et al., 2023) also demonstrate generalization to $\mathcal{D}_{\mathrm{f}}^{\prime}$, compared to simpler MU methods like FT, RL, and GA, which show lower transferability, as evidenced in the class-wise forgetting on (CIFAR-10, ResNet-18).
Table 3: Compositional unlearning on CIFAR-10 and ImageNet-10, where FV represents the forget vector in the original random-data forgetting scenario, and CU-FV denotes compositional unlearning achieved via pre-learned class-wise forget vector arithmetic.
@@
CIFAR-10
Retrain $5.50_{\pm{0.16}}$($0.00$) $11.57_{\pm{0.47}}$($0.00$) $99.88_{\pm{0.05}}$($0.00$) $94.24_{\pm{0.19}}$($0.00$) $0.00$
FV $2.61_{\pm{0.49}}$($2.89$) $8.26_{\pm{1.17}}$($3.00$) $97.33_{\pm{0.47}}$($2.55$) $90.97_{\pm{0.38}}$($3.27$) $2.92$
CU-FV $5.36_{\pm{0.60}}$($0.14$) $9.76_{\pm{0.91}}$($1.81$) $94.93_{\pm{0.64}}$($4.95$) $88.60_{\pm{0.59}}$($5.64$) $3.16$
ImageNet-10
Retrain $4.05_{\pm{0.45}}$($0.00$) $6.60_{\pm{1.07}}$($0.00$) $99.48_{\pm{0.07}}$($0.00$) $96.33_{\pm{0.38}}$($0.00$) $0.00$
FV $2.27_{\pm{0.50}}$($1.78$) $6.13_{\pm{1.40}}$($0.47$) $98.29_{\pm{0.32}}$($1.19$) $95.82_{\pm{0.48}}$($0.51$) $0.99$
CU-FV $2.27_{\pm{1.18}}$($0.00$) $4.95_{\pm{1.86}}$($1.18$) $97.93_{\pm{1.09}}$($0.36$) $91.41_{\pm{1.25}}$($4.41$) $1.49$
@@
Next, we demonstrate the effectiveness of compositional unlearning via forget vector arithmetic (termed CU-FV). Using class-wise forget vectors, we apply their linear combination as defined in (7) to achieve random data forgetting. Tab. 3 compares the performance of CU-FV with Retrain (the exact unlearning method) and the direct forget vector approach (FV) applied to the targeted forget set. Interestingly, we observe that CU-FV achieves the overall performance comparable to FV, as indicated by similar Avg. Gap values. Unlike FV, CU-FV optimizes only the class-wise coefficients in (7), resulting in a much smaller optimization space than FV. However, from UA and MIA-Efficacy metrics, we find that unlearning effectiveness is easier to maintain since unlearning typically targets a smaller subset of data points. In contrast, model utility (RA and TA) may decline more with CU-FV than with FV. This is possibly because retention across a larger set of data points unrelated to unlearning requires a higher level of optimization to preserve model utility effectively.
@@
Assessing Forget Vector via An Input Saliency Lens.
In Figs. 4 and 5, we explore the impact of the forget vector on unlearning and utility retention through input saliency map, using Grad-CAM (Gradient-weighted Class Activation Mapping) (Bengio et al., 2013). Using Grad-CAM, we visualize the salient pixels (i.e., regions most influential to model prediction) for both forget and retain images under different unlearning scenarios: (1) the original model (without unlearning), (2) Retrain, (3) SCRUB-based unlearning, and (4) forget vector-based unlearning. In the first three scenarios, we obtain the input saliency maps on raw forget/retain images without the addition of the forget vector, while in the last scenario, the input saliency is applied to images perturbed by the forget vector. As seen in Fig. 4, only the forget vector and Retrain methods effectively unlearn the target input images, evident from the significant shifts in saliency regions compared to those in the original model. In contrast, the MU baseline SCRUB shows minimal saliency shifts, failing to adequately forget the target data in the last two rows. On the other hand, when evaluating input saliency on retain data in Fig. 5, we observe that the forget vector and other unlearning methods produce input saliency maps similar to those of the original model, indicating retained saliency for non-forgotten images. Moreover, we provide additional visualization results in Appendix D.
@@
In Appendix C, we provide additional ablation studies on the sensitivity of the unlearning-retaining regularization parameter $\lambda_{1}$ in (6) and compare the runtime costs of forget vector calculation with those of other model-based unlearning methods.
@@ 7 Conclusion and Limitations
In this paper, we introduced a novel, data-based approach to machine unlearning (MU) in image classification, termed the forget vector. Unlike traditional model-based MU methods that require retraining or fine-tuning model weights, our approach shows that input-agnostic data perturbations can effectively achieve unlearning objectives. Our method demonstrates competitive performance relative to model-based approximate unlearning techniques. Furthermore, we showcase the potential of compositional unlearning: new forget vectors for unseen tasks, such as unlearning arbitrary subsets across classes, can be generated through simple arithmetic operations, like linear combinations of class-specific forget vectors. Extensive experiments confirm the effectiveness and adaptability of our optimized forget vector.
While effective, the forget vector approach relies on input perturbations, which may not be fully robust against white-box adversarial attacks that are aware of the forget vector strategy. Additionally, compositional unlearning’s dependence on pre-trained class-wise forget vectors could increase computational costs in cases where obtaining these vectors is challenging. Future work could focus on improving the robustness of forget vectors and developing more computationally efficient arithmetic techniques for compositional unlearning.
@@
Appendix A Generalization of MU to Forget Data Shifts

### v4

@@ 6.2 Experiment Results
In table 1, we compare the performance of our forget vector approach with other model-based MU methods across the metrics: UA, RA, TA, MIA-Efficacy, and Avg. Gap vs. Retrain. We highlight two key observations below. First, in terms of unlearning effectiveness (UA and MIA-Efficacy), the data perturbation-based forget vector demonstrates highly competitive performance compared to model update-based MU baselines, mostly ranking among the top two methods with the smallest performance gap relative to Retrain (as evidenced by Avg. Gap). The advantage of the forget vector is particularly evident in MIA-Efficacy, where it usually achieves the closest results to Retrain. Second, in terms of model utility post-unlearning (RA and TA), the forget vector generally leads to a larger performance drop than other methods. This is not surprising, as the forget vector is achieved through data perturbations. However, considering the gain in unlearning effectiveness, the Avg. Gap with Retrain shows that the forget vector remains competitive, ranking among the top two unlearning methods. Third, unlearning methods (including Retrain) do not exhibit the same level of distinctiveness in random data forgetting as it does in class-wise forgetting. This is because in random data forgetting, the retain data could have sufficiently represented the distribution of the forget data, making it more challenging for MIA to distinguish forgotten samples from retained ones. Besides, the corresponding results of various MU methods on ImageNet-10 using VGG-16 can be found in Appendix D.
@@
Table 2: UA (%) of forget vector when transferred to unseen forget sets curated under 3 scenarios on (CIFAR-10, ResNet-18). he results are presented in the same format as Table 1.
@@
Conventionally, unlearning effectiveness is typically measured on the original forget set ($\mathcal{D}_{\mathrm{f}}$). However, with the data perturbation-based forget vector, it is also interesting to investigate its unlearning transferability when applied to a new, previously unseen forget set (denoted as $\mathcal{D_{\mathrm{f}}}^{\prime}$) that share similarities with $\mathcal{D}_{\mathrm{f}}$ and are equally appropriate for unlearning. In the context of class-wise forgetting, we consider $\mathcal{D}_{\mathrm{f}}^{\prime}$ using the testing data from the class targeted for unlearning, where unlearning performance should align closely with Retrain since the test-time data to forget share the same distribution with the training set. In the context of data forgetting, we obtain $\mathcal{D}_{\mathrm{f}}^{\prime}$ by applying the data corruption operation GN1 and ET1 used in fig. 2 to perturb $\mathcal{D}_{\mathrm{f}}$ (last two columns of table 2), where Retrain is no longer the gold standard as training data distribution excludes these shifts, allowing higher UA for better unlearning. As observed in table 2, the unlearning performance of the forget vector remains effective when evaluated on $\mathcal{D}_{\mathrm{f}}^{\prime}$. Among the model update-based unlearning baselines, current SOTA methods such as SCRUB (Kurmanji et al., 2024b), SalUn (Fan et al., 2023) and SSD (Foster et al., 2024a) also demonstrate generalization to $\mathcal{D}_{\mathrm{f}}^{\prime}$, compared to simpler MU methods like FT, RL, and GA, which show lower transferability.
@@
CIFAR-10 ResNet-18 Retrain $5.50_{\pm{0.16}}$($0.00$) $11.57_{\pm{0.47}}$($0.00$) $94.24_{\pm{0.19}}$($0.00$) $99.88_{\pm{0.05}}$($0.00$) $\mathbf{0.00}$
FV $2.61_{\pm{0.49}}$($2.89$) $8.26_{\pm{1.17}}$($3.00$) $90.97_{\pm{0.38}}$($3.27$) $97.33_{\pm{0.47}}$($2.55$) $\mathbf{2.92}$
CU-FV $5.36_{\pm{0.60}}$($0.14$) $9.76_{\pm{0.91}}$($1.81$) $88.60_{\pm{0.59}}$($5.64$) $94.93_{\pm{0.64}}$($4.95$) $\mathbf{3.16}$
ImageNet-10 VGG-16 Retrain $4.05_{\pm{0.45}}$($0.00$) $6.60_{\pm{1.07}}$($0.00$) $96.33_{\pm{0.38}}$($0.00$) $99.48_{\pm{0.07}}$($0.00$) $\mathbf{0.00}$
FV $2.27_{\pm{0.50}}$($1.78$) $6.13_{\pm{1.40}}$($0.47$) $95.82_{\pm{0.48}}$($0.51$) $98.29_{\pm{0.32}}$($1.19$) $\mathbf{0.99}$
CU-FV $2.27_{\pm{1.18}}$($1.78$) $4.95_{\pm{1.86}}$($1.65$) $91.41_{\pm{1.25}}$($4.92$) $97.93_{\pm{1.09}}$($1.55$) $\mathbf{2.47}$
ImageNet-10 ViT Retrain $1.41_{\pm{0.06}}$($0.00$) $99.07_{\pm{0.01}}$($0.00$) $93.57_{\pm{0.00}}$($0.00$) $99.27_{\pm{0.01}}$($0.00$) $\mathbf{0.00}$
FV $1.08_{\pm{0.39}}$($0.33$) $98.97_{\pm{0.35}}$($0.10$) $91.40_{\pm{1.30}}$($2.17$) $99.10_{\pm{0.50}}$($0.17$) $\mathbf{0.69}$
CU-FV $1.62_{\pm{0.05}}$($0.21$) $98.55_{\pm{0.14}}$($0.52$) $92.09_{\pm{0.25}}$($1.48$) $98.50_{\pm{0.01}}$($0.77$) $\mathbf{0.75}$
Table 3: Compositional unlearning on CIFAR-10 and ImageNet-10 for random data forgetting, where FV represents the original setting of forget vector that is directly learned based on a targeted forget set, and CU-FV denotes compositional unlearning achieved via pre-learned class-wise forget vector arithmetic. The results are presented in the same format as Table 1.
@@
Next, we demonstrate the effectiveness of compositional unlearning via forget vector arithmetic (termed CU-FV). Given pre-computed class-wise forget vectors, we apply their linear combination as defined in (7) to achieve random data forgetting. table 3 compares the performance of CU-FV with Retrain (the exact unlearning method) and the direct forget vector approach (FV) applied to the targeted forget set. Interestingly, we observe that CU-FV achieves the overall performance comparable to FV, as indicated by similar Avg. Gap values. Unlike FV, CU-FV optimizes only the class-wise coefficients in (7), resulting in a much smaller optimization space than FV. However, from UA and MIA-Efficacy metrics, we find that unlearning effectiveness is easier to maintain since unlearning typically targets a smaller subset of data points. In contrast, model utility (RA and TA) may decline more with CU-FV than with FV.
Assessing Forget Vector via An Input Saliency Lens.
In figs. 4 and 5, we explore the impact of the forget vector on unlearning and utility retention through input saliency map, using Grad-CAM (Gradient-weighted Class Activation Mapping) (Bengio et al., 2013). Using Grad-CAM, we visualize the salient pixels (i.e., regions most influential to model prediction) for both forget and retain images under different unlearning scenarios: (1) the original model (without unlearning), (2) Retrain, (3) SCRUB-based unlearning, and (4) forget vector-based unlearning. In the first three scenarios, we obtain the input saliency maps on raw forget/retain images without the addition of the forget vector, while in the last scenario, the input saliency is applied to images perturbed by the forget vector. As seen in fig. 4 the forget vector and Retrain effectively unlearn the target input images, evident from the significant shifts in saliency regions compared to those in the original model. In contrast, the MU baseline SCRUB shows minimal saliency shifts, failing to adequately forget the target data in the last row. On the other hand, when evaluating input saliency on retain data in fig. 5, we observe that the forget vector and other unlearning methods produce input saliency maps similar to those of the original model, indicating retained saliency for non-forgotten images. Additional visualization results can be found in Appendix G.
@@
In Appendix E, we provide additional ablation studies on the sensitivity of the unlearning-retaining regularization parameter $\lambda_{1}$ in (6). Moreover, we perform the efficiency analysis of forget vector calculation with those of other model-based unlearning methods in Appendix F.
@@ 7 Conclusion and Limitations
In this paper, we introduce a novel, data-based approach to machine unlearning (MU) in image classification, termed the forget vector. Unlike traditional model-based MU methods that require retraining or fine-tuning model weights, our approach shows that input-agnostic data perturbations can effectively achieve unlearning objectives. Our method demonstrates competitive performance relative to model-based approximate unlearning techniques. Furthermore, we showcase the potential of compositional unlearning: new forget vectors for unseen tasks, such as unlearning arbitrary subsets across classes, can be generated through simple arithmetic operations, like linear combinations of class-specific forget vectors. Extensive experiments confirm the effectiveness and adaptability of our optimized forget vector.
While effective, the forget vector approach relies on input perturbations, which may not be fully robust against white-box adversarial attacks that are aware of the forget vector strategy. Additionally, compositional unlearning’s dependence on pre-trained class-wise forget vectors could increase computational costs in cases where obtaining these vectors is challenging. Future work could focus on improving the robustness of forget vectors and provide theoretical guarantees for compositional unlearning.
@@
Appendix A Details of MIA Implementation
To evaluate the effectiveness of the unlearning process, MIA is implemented following (Jia et al., 2023) using a prediction confidence-based attack method (Song & Mittal, 2021), which comprises a training phase and a testing phase in its computation. Specifically, a balanced dataset is first formed by sampling data points from retain set $\mathcal{D}_{\mathrm{r}}$ and test set $\mathcal{D_{\mathrm{t}}}$, ensuring the distinction from the forget set $\mathcal{D_{\mathrm{f}}}$. Then, a MIA predictor is trained utilizing such balanced dataset. Thereafter, MIA-Efficacy can be calculated by applying the trained MIA predictor to the unlearned model ${{\bm{\theta}}_{\mathrm{u}}}$ on the forget set $\mathcal{D_{\mathrm{f}}}$. Essentially, the goal is to determine how many samples in $\mathcal{D_{\mathrm{f}}}$ can be accurately identified as non-training data with respect to ${{\bm{\theta}}_{\mathrm{u}}}$ by the MIA model. Formally, MIA-Efficacy is defined as follows,
[…]
where $N_{tn}$ represents the total number of true negative samples in the forget set $\mathcal{D_{\mathrm{f}}}$ predicted by the trained MIA model, i.e., the number of forgetting samples classified as non-training examples.
Appendix B Generalization of MU to Forget Data Shifts